Ali is twice as old as Ed. Four years ago, he was four times as old as ed. When will the sum of their ages be 66?
Answer by alicealc(293)

Ali's age = A
Ed's age = E

A = 2E

A - 4 = 4*(E -4)
2E - 4 = 4E - 16
16 - 4 = 4E - 2E
12 = 2E
E = 12/2 = 6

A = 2E = 2*6 = 12

6 + x + 12 + x = 66
18 + 2x = 66
2x = 66 - 18
2x = 48
x = 48/2 = 24

so 24 years later the sum of their ages will be 66